OpenStudy (anonymous):

the area of a rectangular field is x^2-x-72 m^2. the length of the field is x+8 m(meters). what is the width of the field in meters?

OpenStudy (anonymous):

You know that area=(length)(width) So you have: \[x^2-x-72=(x+8)*w\] Solve for w.

OpenStudy (anonymous):

wow i appreciate it, thank you for your quick answer

OpenStudy (anonymous):

so W would be?

OpenStudy (anonymous):

the length would be 8?

OpenStudy (anonymous):

Well the left side factors to: \[(x-9)(x+8)=(x+8)w \implies w=(x-9)\]

OpenStudy (anonymous):

i find trouble resolving this clearly arg

OpenStudy (anonymous):

So w would be x-9 meters

OpenStudy (anonymous):

yes but the x would be?

OpenStudy (anonymous):

or there is no clear answer because it doesnt tell right ?

OpenStudy (anonymous):

You can't solve it from the information given. You know l in terms of x, you know w in terms of x, and you know area in terms of x, you can't solve x explicitly the way I see it.

OpenStudy (anonymous):

so there is no term since it is not given we finish at W=x-9 thank you very much for your patience

OpenStudy (anonymous):

No problem